OpenStudy (anonymous):

What is the solution to the system of equations? 3x-4y=19 y=2x-11 a. (5,-1) b. (-5,-21) c. (5,1) d. (-5,-8.5) please show work! also please tell the method you used to solve it! thanks!

OpenStudy (anonymous):

ok put 2x-11 in for y in the top equation to start with. 3x-4(2x-11)=19

OpenStudy (anonymous):

distribute the 4 3x-8x-44=19 -5x-44=19

OpenStudy (anonymous):

solve for x -5x-44=19 -5x=25 x=-5

OpenStudy (anonymous):

then substitute -5 for x in your second equation and solve y= 2(-5)-11 y=-10-11 y=-21 and that's it